# PTMPicker: Facilitating Efficient Pretrained Model Selection for Application Developers

### v1

@@ 5.4.1. RQ1: Reliability of metadata extraction
However, automatically assessing these aspects is challenging due to the absence of ground truth, thereby necessitating manual verification. Moreover, it is impractical for us to manually inspect the metadata extraction from all 5,069 experimental PTMs, as this would be excessively time-consuming. Therefore, we randomly sampled and manually evaluated 100 models from the curated PTM collection, referencing the evaluation approach in (Jiang et al., 2024).
@@
Finally, two independent authors achieved high inter-rater reliability, as measured by Cohen’s unweighted kappa coefficient with the value of 0.81 (coh, 2024), which means almost perfect agreement between two raters. As a result of this process, the accuracy of metadata extraction based on the manually inspected sample of 100 model cards reached 95%. To complement and automate this manual inspection, we enhanced the validation process by crafting specialized prompts with few-shot examples to assess the validity of the metadata extraction. These prompts were designed to simulate the human validation process, in which the correctness of extracted metadata is assessed against the corresponding model card. Specifically, the few-shot examples embedded in these prompts guided the language model to more accurately assess the validity of each extracted attribute. Ultimately, this prompt-based validation approach achieved an accuracy of 95%, matching the manual inspection results and demonstrating its effectiveness as a scalable alternative.
@@
5.4.2. RQ2: Characteristics of metadata
Fig.5 presents the number of models associated with each summarized metadata attribute. The three most frequently provided metadata fields among the selected PTMs are licensefunction, and dataset. Among these, the function is extracted from the model’s associated model card using GPT-4 with carefully crafted prompts, while the others can be directly specified by model developers through the Hugging Face metadata UI or by editing the associated YAML file(Hugging Face, ndb) and thus are straightforwardly scraped from the model hosting site. In contrast, most of the remaining metadata fields are not explicitly required by the hosting platform, and thus can only be extracted from the free-form content of the model card. This suggests that, in the absence of mandatory fields, authors are generally less inclined to voluntarily provide detailed metadata.
Fig. 7 and Fig. 7 present the average and median downloads for each extracted metadata field, depending on whether the metadata field is introduced in the model card. In general, PTMs associated with certain metadata tend to exhibit higher downloads, a trend also observed in prior work (Jones et al., 2024). This correlation suggests that the presence of certain metadata improves the visibility, perceived quality, or usability of PTMs, thereby contributing to their increased adoptions. However, certain metadata fields, including basemodel, framework, software, copyright, modelsize, inference cost, and carbon emitted, exhibit an opposite trend, where PTMs associated with these attributes tend to have fewer downloads. One possible explanation is that users may perceive these metadata fields as less critical or less directly relevant to their application needs. For instance, metadata such as the model’s base model (i.e., the upstream model used for pretraining or fine-tuning) may be regarded as implementation-specific and of limited relevance to functional performance. Additionally, the infrequent provision of these specific metadata may contribute to their lower perceived importance and, consequently, to fewer downloads.
@@ 5.4.4. RQ4: Accuracy of model selection
Figure 8. Evaluation results for pretrained model selection.

### v3

@@ 5.4.1. RQ1: Reliability of metadata extraction
However, automatically assessing these aspects is challenging due to the absence of ground truth, thereby necessitating manual verification. Moreover, it is impractical for us to manually inspect the metadata extraction from all 5,069 experimental PTMs, as this would be excessively time-consuming. Therefore, we randomly sampled and manually evaluated 100 models from the curated PTM collection, referencing the evaluation approach in (jiang2024peatmoss).
@@
Finally, two independent authors achieved high inter-rater reliability, as measured by Cohen’s unweighted kappa coefficient with the value of 0.81 (cohenkappa), which means almost perfect agreement between two raters. As a result of this process, the accuracy of metadata extraction based on the manually inspected sample of 100 model cards reached 95%. To complement and automate this manual inspection, we enhanced the validation process by crafting specialized prompts with few-shot examples to assess the validity of the metadata extraction. These prompts were designed to simulate the human validation process, in which the correctness of extracted metadata is assessed against the corresponding model card. Specifically, the few-shot examples embedded in these prompts guided the language model to more accurately assess the validity of each extracted attribute. Ultimately, this prompt-based validation approach achieved an accuracy of 95%, matching the manual inspection results and demonstrating its effectiveness as a scalable alternative.
@@ 5.4.2. RQ2: Characteristics of metadata
Fig.5 presents the number of models associated with each summarized metadata attribute. The three most frequently provided metadata fields among the selected PTMs are licensefunction, and dataset. Among these, the function is extracted from the model’s associated model card using GPT-4 with carefully crafted prompts, while the others can be directly specified by model developers through the Hugging Face metadata UI or by editing the associated YAML file(modelmd:2025) and thus are straightforwardly scraped from the model hosting site. In contrast, most of the remaining metadata fields are not explicitly required by the hosting platform, and thus can only be extracted from the free-form content of the model card. This suggests that, in the absence of mandatory fields, authors are generally less inclined to voluntarily provide detailed metadata.
Fig. 7 and Fig. 7 present the average and median downloads for each extracted metadata field, depending on whether the metadata field is introduced in the model card. In general, PTMs associated with certain metadata tend to exhibit higher downloads, a trend also observed in prior work (jones2024we). This correlation suggests that the presence of certain metadata improves the visibility, perceived quality, or usability of PTMs, thereby contributing to their increased adoptions. However, certain metadata fields, including basemodel, framework, software, copyright, modelsize, inference cost, and carbon emitted, exhibit an opposite trend, where PTMs associated with these attributes tend to have fewer downloads. One possible explanation is that users may perceive these metadata fields as less critical or less directly relevant to their application needs. For instance, metadata such as the model’s base model (i.e., the upstream model used for pretraining or fine-tuning) may be regarded as implementation-specific and of limited relevance to functional performance. Additionally, the infrequent provision of these specific metadata may contribute to their lower perceived importance and, consequently, to fewer downloads.
@@ 5.4.4. RQ4: Accuracy of model selection
Figure 8. Evaluation results for pretrained model selection.